OpenStudy (anonymous):

What is the 185th digit in the following pattern 12345678910111213141516.... (A) 0 (B) 5 (C) 7 (D) 8 (E) 9

OpenStudy (anonymous):

@kelliegirl33 @nincompoop @UditKulka

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Can you explain on how you approached your answer?

OpenStudy (anonymous):

yes. 1 to 9: 9 digits 10-99: 200 digits 9+200 is 24 more than 185, so the last number is 12 behind 99. 99-12=87 the pattern goes up to ...858687, and 7 is the 185th digit.

OpenStudy (anonymous):

1 to 9: 9 digits 10-99: 200 digits 9+200 is 24 more than 185, so the last number is 12 behind 99. 99-12=87 the pattern goes up to ...858687, and 7 is the 185th digit

OpenStudy (anonymous):

Did you get this from Yahoo Answers? @kayla_b_stephens @UditKulka

OpenStudy (anonymous):

1 digit numbers 1 to 9 used = 9 total digits = 9 2 digit numbers 10 to 99 used = 90 total digits = 180 if numbers are used from 1 to 99 digits used = 189 you require 185 digits so cancell 4 digits from the end 123...959697(9899) so required digit = 7 other wise 1 digits used = 9 digits remained = 185 – 9 = 176 176 digits are by using 2 digits so numbers used = 176/2 = 88 10, 11, T88 = 10 + 87 = 97 required digit = 7 ----

OpenStudy (anonymous):

yes but it is true

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

no problem